Expanded form of 89 274 069


1
Expert's answer
2021-03-30T05:45:38-0400

"9\times 1=9"

"6\times10=60"

"0\times 100=0"

"4\times1000=4000"

"7\times 10000=70000"

"2\times100000=200000"

"9\times1000000=9000000"

"8\times10000000=80000000"


"\therefore 80,000,000+9,000,000+200,000+70,000+4,000+00+60+9" is the expanded form of 89274069